Qualitative behavior of Volterra difference systems

J. Morchało, Poznań, Poland

Keywords: Volterra difference systems, asymptotic behavior, ρ-admissible maps AMS Subject classification: 39A10, 39A12, 39A70

Abstract: In this paper we study the behavior of the solutions of Volterra difference system of the form (N) treated as a perturbation of the linear system (L) or (E). We shown that (L) inherits much of the behavior (E) for an appropriate perturbation term C(n)x.